OpenStudy (anonymous):

Evaluate the following expression. 2^-3

OpenStudy (usukidoll):

hey! Long time no see ! :D

OpenStudy (anonymous):

Hi!

OpenStudy (usukidoll):

so we are given the problem \[\LARGE 2^{-3} \] negative exponents aren't allowed, so we have to take the reciprocal. A negative exponent just means that the base is on the wrong side of the fraction line, so you need to flip the base to the other side. For example \[ \LARGE x^{-2} \rightarrow \frac{1}{x^2} \]

OpenStudy (usukidoll):

so we need to flip \[\LARGE 2^{-3} \] and that becomes ..?

OpenStudy (anonymous):

-3 over 2?

OpenStudy (usukidoll):

not exactly... remember our example \[\LARGE x^{-2} \rightarrow \frac{1}{x^2} \] so let x =2 and replace the -2 with -3

OpenStudy (usukidoll):

A negative exponent is equivalent to the inverse of the same number with a positive exponent

OpenStudy (usukidoll):

example \[\LARGE x^{-4} \rightarrow \frac{1}{x^4} \]

OpenStudy (anonymous):

Oh okay I got it!

OpenStudy (usukidoll):

so let's try to apply the example to \[\LARGE 2^{-3} \]

OpenStudy (usukidoll):

so instead of \[\LARGE 2^{-3} \] we have \[\LARGE 2^{-3} \rightarrow \frac{1}{?} \]